OpenStudy (anonymous):

In an examination, a student who got 20% of the total marks fails by 18 marks. Had he got 32% of the total marks he would have got 6 marks more than passing marks. What is the passing marks?

OpenStudy (lgbasallote):

which a will be first....

OpenStudy (anonymous):

32-12 percent of the full marks=12 percent of the full marks=18+6=24 marks so, full marks= 24/12 *100=200

OpenStudy (lgbasallote):

seems arnab got first :p

OpenStudy (anonymous):

more important, @Pratika_Rathore , u got it?

OpenStudy (kropot72):

Pass mark is 58

OpenStudy (apoorvk):

So let the maximum marks be x, and the passing marks be y. he gets 20% of it - that is one fifth x. which x/5 now if he had scored 18 more than x/5, more his score would have reached passing score, which is 'y' so, x/5 + 18 = y ...{i} If he scored 32% of the full marks, that (32/100)*x, his score would have been '6' more than the passing score 'y' so, 32 ---x = y + 6 ....{ii} 100 Two linear equations 'i' and 'ii' - solve them to find out 'x' and 'y', and enjoy!! Post your answer to verify.

OpenStudy (anonymous):

@arnab but no such option is mentioned

OpenStudy (anonymous):

yeah, i left u to do the rest

OpenStudy (anonymous):

Ya got it@apoorvk x=200 and y=58 thanks bro

OpenStudy (kropot72):

Let t = total marks 0.32t - 6 = pass mark 0.2t + 18 = pass mark Solving for t gives t = 200 Substituting for t gives pass mark = 58

OpenStudy (apoorvk):

lol, you're (right + welcome). :P